OpenStudy (anonymous):

Let f(x)=ax^3+bx^2+cx+d and suppose that r is a root of the equation f(x)=0. show that -r is a root of the equation f(-x)=0

OpenStudy (anonymous):

you know that: ar^3+br^2+cr+d=0 now look at f(-x) = -ax^3+bx^2-cx+d plug x=-r -a*(-r)^3+b*(-r)^2-c*(-r)+d = ar^3+br^2+cr+d and this is 0 according to the second line :)

OpenStudy (anonymous):

my "proof" is f(r)=0 and if -x=-r then f(-x)=f(-r)=0

OpenStudy (anonymous):

does that work

ganeshie8 (ganeshie8):

how did u get f(-x) = f(-r) = 0 ? @muzzammil.raza you need to prove it using coolsector work first right ?

OpenStudy (anonymous):

ok thanks

ganeshie8 (ganeshie8):

:)